OpenStudy (samara8954):

The Mendez family used 12 gallons of gasoline on a 300-mile trip. At that rate, how far would they go on 9 gallons of gasoline?

OpenStudy (samara8954):

do you divide

OpenStudy (anonymous):

Use proportions 2 Gallons / 300 Mile trip \[ \frac{ 12 ~Gallons}{300 ~ Miles}=\frac{9 ~Gallons}{x ~Miles}\] \[\frac{12}{300}=\frac{9}{x}\] Cross multiply and solve for x

OpenStudy (anonymous):

12*x=9(300) 12x=2700 Now divide 12 from both sides and we get x=225

OpenStudy (samara8954):

so the answer is 25

OpenStudy (anonymous):

225

OpenStudy (anonymous):

oh gosh -.-

OpenStudy (anonymous):

my golly gosh

OpenStudy (anonymous):

if for 12 gallons we drove 300 miles then for 9 gallons we drove for 225

OpenStudy (samara8954):

so the answer is 225

OpenStudy (samara8954):

not 25

OpenStudy (samara8954):

what?

OpenStudy (anonymous):

yup yup

OpenStudy (samara8954):

A restaurant staff made 54 sandwiches for 18 customers. At that rate, how many sandwiches would the restaurant staff make for 25 customers?

OpenStudy (anonymous):

THE ANSWER IS 225 :D

OpenStudy (samara8954):

A. 75 B. 79 C. 43 D. 72

OpenStudy (samara8954):

help me with the new question

OpenStudy (samara8954):

what is the answer